if X and Y are the differnent roots of the equation acosA+ bsinA=c then cos[X+Y]=....

Grade:11

1 Answers

aditi arora
8 years ago
AS  X and Y are the differnent roots of the equation acosA+ bsinA=c 
a cos A = c- bsin A 
squaring both sides
a2 ( cos A )2 = c+ b2sin2A
cos2A = i – sin2A
(a2+b2) sin2A – 2acsin A + c2-b2 = 0
sinXsinY= c2-b2 / a2+b2                                                     similarly solve for cosX cos Y = c2-a/ a2+b
cos (X+Y) = cos X cos Y – sin X sin Y
PUT VALUES and you’ll get answer